OpenStudy (anonymous):

HHHEEEELLLPPPP INVERSES f(x)e^2x

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

hmm have you covered logarithms yet?

OpenStudy (jdoe0001):

I assume is \(\large \bf f(x)=e^{2x}\) you'd end up with a logarithmic function though, thus... have you covered logarithms yet?

OpenStudy (anonymous):

would you take the natural log?

OpenStudy (anonymous):

and yes

OpenStudy (anonymous):

\[\frac{ x }{ e^2 }\]

OpenStudy (anonymous):

Maybe?

OpenStudy (anonymous):

\[\ln x =\ln e ^{2y}\]

OpenStudy (anonymous):

then move the 2y in front?

OpenStudy (jdoe0001):

ok.... so... let us use the log cancellation rule of . actually... yes... let us use the exponent instead

OpenStudy (anonymous):

and In e = 1 so it would be...\[\ln x=2y\]

OpenStudy (anonymous):

oh okay

OpenStudy (jdoe0001):

\(\bf f(x)=y=e^{2x}\qquad inverse\implies x=e^{2y} \\ \quad \\ log_{\color{brown}{ e}}x=log_{\color{brown}{ e}}({\color{brown}{ e}}^{2y})\implies log_e(x)=2y{\color{purple}{ log_e(e)}}\implies ln(x)=2y\cdot {\color{purple}{ 1}} \\ \quad \\ \cfrac{ln(x)}{2}=y\)